Estimate how many collisions a water molecule undergoes at room temperature in 1 μs. Envision a liquid at a temperature of 1 K, where the distance between molecules and the molecular mass are the same as those of water at room temperature? Estimate how many collisions a molecule therein undergoes in 1 μs (at a temperature of 1 K).
Answer #1

Collision frequency formula:

R= universal gas constant = 8.314 J K-1 mol-1

T= room temperature = 25 degree Celsius = 25×273 =298 K

M= molar mass= 1 kg mol-1

= 4.99 m s-1

~ 5 m s-1

R = 8.314 × 10-5 bar m3 K-1mol-1

T = 298 K

= molecule' s effective collision area = 2d

For water d= 0.28 nm

= 2× .28 nm= .56×10-9 M

NA= 6.023×1023 mol-1

P = .317 bar at room temperature

=( 2477.57 ×10-5 m)/(2.658×105 )

= 93.2 nm

= (5 m s-1 )/93.2 nm

= 53.6 ×106 s-1 =(53.6×106)/ 1s

=53.6 × 106 ×10-6

Number of collisions per micro second = 53.6